OpenStudy (anonymous):

derivative f(x)=3^(4x^5+x-8)*log[2](6x+3)

myininaya (myininaya):

\[y=3^{4x^5+x-8} \cdot \log_2(6x+3) ?\]

OpenStudy (anonymous):

yep thats the one!

myininaya (myininaya):

so we have a product so we need the product rule y'=f'g+fg' right?

myininaya (myininaya):

Now lets look at f \[f=3^{4x^5+x-8} ; g=\log_2(6x+3)\]

myininaya (myininaya):

oops lets look at f first ok?

myininaya (myininaya):

Do you know how to find derivative of that?

OpenStudy (anonymous):

no

myininaya (myininaya):

\[f=3^{p(x)}\] to find derivative of this take ln( ) of both sides

myininaya (myininaya):

\[\ln(f)=\ln(3^{p(x)})\]

myininaya (myininaya):

\[\ln(f)=p(x) \ln(3)\]

myininaya (myininaya):

Take derivative of both sides

OpenStudy (anonymous):

soryy but could u please explane me a little bit more! i am really bad in derivatives!

myininaya (myininaya):

f and p are both functions of x

myininaya (myininaya):

(ln(f))'=?

myininaya (myininaya):

f'/f right?

OpenStudy (anonymous):

f=3^(4x^5+x−8)'/3^(4x^5+x−8) thats right

myininaya (myininaya):

no no

OpenStudy (anonymous):

i see i am really bad at this :)

myininaya (myininaya):

\[\frac{f'}{f}=\ln(3) p'(x)\] <---i'm trying to find f'

myininaya (myininaya):

we need to solve this for f'

myininaya (myininaya):

how do we do it?

OpenStudy (anonymous):

no idea! :) thats the one i dont know how to do!

myininaya (myininaya):

you can solve for f' just multiply f on both sides and therefore f' is isolated

OpenStudy (anonymous):

\[\frac{d}{dx}b^x=b^x\ln(b)\] so by the chain rule \[\frac{d}{dx}3^{4x^5+x-8}=\ln(3) 3^{4x^5+x-8}\times (20x+1)\]

OpenStudy (anonymous):

\[\frac{d}{dx}\log_b(x)=\frac{1}{x\ln(b)}\]so again by the chain rule \[\frac{d}{dx}\log_2(6x+3)=\frac{1}{(6x+3)\ln(2)}\times 6\]

OpenStudy (anonymous):

clean it up to get \[\frac{d}{dx}\log_2(6x+3)=\frac{3}{(2x+1)\ln(2)}\] and then use the product rule to get your answer

OpenStudy (anonymous):

\[(fg)'=f'g+gf'\] plug in the functions and the derivatives, don't simplify

OpenStudy (anonymous):

thank u!